OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}=\frac{4-3}{4-1}=\frac{1}{3}\] \[y-4=\frac{1}{3}(x-4)\] \[y-4=\frac{1}{3}x-\frac{4}{3}\] \[y=\frac{1}{x}x+\frac{8}{3}\]

OpenStudy (anonymous):

First find the slope of the two equations by plugging them in: (y2 -y1)/ (x2-x1) (1,3) = (x1, y1) (4, 4) = (x2, y2) Solve and you should the slope or m in the equation y = mx + b Then plug the "m" value in the equation: y - y1) = m (x - x1) This time, pick any of the two coordinates provided as being (x1, y1) Doesnt matter which one.. Solve and you should get your standard equation =)

OpenStudy (anonymous):

so the equation is y-3=1/3 (x-1) then how would you solve that? I know you have to add 3 to both sides first

OpenStudy (anonymous):

y-y1/y2-y1=x-x1/x2-x1 y-3/4-3=x-1/4-1 y-3/1=x-1/3 3(y-3)=1(x-1) 3y-9=x-1 x-3y+9-1 x-3y+8=0 this is the eqaation of two points

OpenStudy (anonymous):

okay i got it now! thanks alot! everythings right execpt its - 8 instead of +8 :)